What is the pH of a 0.20 M solution of (a) H3PO4, (b) KH2PO4, (c) K2HPO4, and (d) K3PO4?

The answers are (a) 1.46, (b) 4.67, (c) 9.76, and (d) 12.74. I have no idea how to do these problems, please show me step by step how to get to these answers.
Answer #1

3 H2 PO 70 1 O 2 0.2ース 15 X 10-32 ズ2 2 17.5X10-3102 2 e g 0:037 KH PO ka= 62x10-82t =|HtJ っ 3.lo×10-7 6:51 PO に一10-1 po 0.2- honed 0.0208×十×2= O -0.oou 17